Use the data below to find the growth of income per person (over the entire period, not an annual basis) between the two years listed.

Year

Real GDP

Population

Income per person

1990

$ 3,460,000 million

245 million

?

2010

$ 6,680,000 million

300 million

?

Income per person grew by (in decimal percent, i.e. .32 not 32).=
Answer #1

Answer

Income per person =real GDP/population

the income per person in 1990=3460000/245=14122.449

the income per person in 2010=6680000/300=22266.6667

Growth =((amount of the year -the amount of last year )/amount of last year)*100

=((22266.6667-14122.449)/14122.449)*100

=57.6685935

the income grew by $57.67%